Prev Next

SW&IT Meeting Minutes 2021-02-01

Attendees: David, Igor, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210125
LR20210125 - Bec, Matthieu 40.0
LR20210125 - Edmeades, David 0.0
LR20210125 - Sola, Igor 0.0
LR20210125 - Grenz, Paul 0.0
LR20210125 - Hooper, Stephen 0.0
LR20210125 - Zhang, Yang 40.0
LR20210125 - Zhang, Xianyu 40.0

News from management

  • Strategic planning recap.

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT S/W RIX responses
    • Stategic planning x2 1/2days
    • time off Mon
  • This week
    • SHARK NIR RFAs
    • AGW OAC encoder update meeting
    • DMS react.js
    • Web student posting

DE

  • Archive stuff; no response from Cristina yet. I'll follow up
  • Server room power shutdown recovery
  • Cobbler repo mirror dive; next is command line work
  • Started fresh with storage doc for Weds
  • Early reminder re: power work 13-15(?) Feb
  • Focus push for Cisco training
  • Planning trip next Monday to replace ailing disk 5; any other tasks let me know

YZ

  • Previous:
    • Some TCS work (AOSGUI patch, draft new releasing doc).
    • TMS work and meeting.
    • OSCO meeting and Cravat/OSCO software work
  • Next:
    • More TCS work (IIF patch on Tuesday, decided to revise the overall TCS releasing/dev doc).
    • Cravat/OSCO work to continue...
    • TMS work to continue...

XZ

  • Last week
    • Answer questions for ATT AdSec and SW
    • AO support/AdSec SW update
    • SOUL/AdSec meetings
  • This week
    • Work on Seveal SX AdSec issues (DSP watchdog expires, HK not receives telemetry, FD telemetry stops)
    • FSM Fault recovery
    • AO support

SH

  • strategic planning meeting
  • Tucson computer room shutdown
  • UITS third phase things

next:

  • core-b work on Tuesday
  • meeting with UITS this week about 10G in Tucson CR
  • Steward meeting on Wednesday
  • DIMM stuff

IS

Last week:
  • Luci AO field measurement
  • LUCI
    • IT #7493 Aladin slow subtraction. Doing the subtraction outside Aladin
    • IT #8231- Timeout on GEIRS during readout
  • PIT
    • Generate new CSV file for archive and for OT

This week:
  • LUCI
    • IT #7493 Aladin slow subtraction
    • IT #8231- Timeout on GEIRS during readout
    • IT #8026- Possibly implementing Update OT with new AO patrol field parameters.
    • create a luci rpm via a docker container (sorry! I keep forgetting)

PSG

  • Last week
    • Testing of swing arm, commands work now.
  • This week
    • Adding code to OSS
Topic revision: r12 - 01 Feb 2021, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback